# Momentum dependent quantum Ruelle-Pollicott resonances in translationally invariant many-body systems

### v2

@@ IV.1 Power-law decay
Figure 10: Eigenvalue spectrum of $M(k=0^{-})$ for $\tau=0.45,h_{\rm x}=0.9,h_{\rm z}=0.8$ and $r=7$. There are $5985$ eigenvalues, with the largest at $\lambda_{1}\approx 0.914$ being highlighted by a red circle, with further 10 with the largest modulus being shown with a small blue triangle (three largest ones are $\lambda_{2}\approx 0.897$, and a complex pair $|\lambda_{3,4}|\approx 0.861$). Blue circles are theoretical $r_{\rm in}$ and $r_{\rm out}$ (18).
@@ IV.2 Prethermalization in $k=0^{+}$
Considering the above explanation we must also rethink the apparent plateau seen in the $k=0^{+}$ sector (Fig. 8). Namely, it seems more natural that one would have a decay in all sectors. And indeed, looking at the autocorrelation function of $Z$ (20) at much longer times we eventually do see a decay, Fig. 11. At a finite $L$ there appears to be a 2nd plateau at a late time, e.g. $t>10^{5}$ at $L=20$, which though, as seen in the figure, is just a finite-size effect. As one increases $L$ one approaches the asymptotic exponential decay given by the largest eigenvalue that is very close to $1$, namely $\lambda_{1}\approx 1-14\cdot 10^{-6}$ (for studied truncations $\lambda_{1}$ seems rather stable with only small even-odd $r$ oscillations). This exponential decay only starts after very long transient plateau of duration $\sim 10^{3}$. Looking at the $\tau$ dependence in the inset of Fig. 11 we see that the time and the height of this 2nd finite-size plateau rapidly change with $\tau$. We therefore have a rather interesting situation: observables from the $k=0^{-}$ sector decay (quickly) as a power law, while those from the $k=0^{+}$ eventually do decay exponentially but only after a long plateau.
The long plateau can be explained by what has been called a prethermalization keiji16 . Namely, although $\tau$ does not necessarily look perturbatively small one can nevertheless try to write the effective one-step Floquet operator $H_{\rm eff}$ using a Baker-Campbell-Hausdorff (BCH) expansion
@@
with $H_{p}$ being the antihermitian $p$-th order term (in $\tau$) in the BCH expansion. If this expansion would converge, and therefore could describe the long-time dynamics by simple $U^{t}={\rm e}^{H_{\rm eff}t}$, the $H_{\rm eff}$ would be a constant of motion. Provided it would be a local operator this should be reflected in the spectral properties of $M(k=0^{+})$, i.e., one eigenvalue being $1$. However, the formal BCH expansion is usually divergent with terms $H_{p}$ eventually growing for $p>p_{0}$. Up until times when divergent $H_{p>p_{0}}$ become important the $H_{\rm eff}$ truncated at the order $p_{0}$ is almost conserved keiji16 , i.e., its exact lack of conservaton will become noticeable only at large times. Such an almost conserved $H_{\rm eff}$, a hallmark of prethermalization keiji16 , will result in one eigenvalue of $M(k=0^{+})$ being very close to $1$. And this is exactly what happens in our kicked Ising model at $\tau=0.45$. This can be seen in Fig. 12. As already mentioned, the largest eigenvalue $\lambda_{1}$ is very close to $1$ with the corresponding eigenvector $\mathbf{x}_{1}$ being, upto normalization, exactly equal to the truncated $H_{\rm eff}$. This is verified in the bottom frame of Fig. 12 where we calculate the overlaps between $\mathbf{x}_{1}$ calculated from $M(0^{+})$ for a finite support size $r$, say $r=3$, and compare it with $H_{\rm eff}$ obtained by summing up the first $p$ orders of the BCH series. Due to the structure of the kicked Ising model the support of the BCH term $H_{p}$ of order $p$ is $\sim 1+p/2$, i.e., every 2nd order increases the support of commutators by $1$ site. This can be nicely seen from the data: for $r=3$ the fidelity saturates at $p\approx 6$, while at $r=5$ at around $p\approx 10$. One can also see that the error of $\mathbf{x}_{1}$ for a given $r$, i.e., its difference to $H_{\rm eff}$, is essentially equal to the norm $|H_{p}|^{2}:={\rm tr}(H_{p}^{\dagger}H_{p})$ of the truncated terms (red diamonds in the figure) for $p\approx 2r$. This all works as long as the BCH series converges, i.e., $H_{p}$ decrease. Namely, by calculating BCH terms upto order $20$ casas08 we can see (Fig. 12) that the terms indeed start to increase after $p_{0}\approx 16$. This means that $H_{\rm eff}$ is not really conserved and we do not expect any eigenvalue of $M(k=0^{+})$ to be exactly $1$. The eventual divergent nature of the BCH series is also reflected in the structure of the eigenvector $\mathbf{x}_{1}$ – for larger $r$ it ceases to be well approximated by $H_{\rm eff}$ (compare $r=5$ and $r=7$ in the figure) – as well as in the spectrum of $M(0^{+})$ (other eigenvalues moving closer to $\lambda_{1}$).
@@
One question that remains is what happens in the thermodynamic limit $L\to\infty$, in particular in view of a seeming power-law decay in other sectors? There are two possibilities. As we can see in the top Fig. 12 there are many eigenvalues on the real axis, the number of which increases with $r$, similarly as is the case in the $k=0^{-}$ sector. So it could be that at very large $r$ when those eigenvalues move really close to $1$ they begin to mix with $\lambda_{1}$ which eventually looses its “prethermalization” character of describing $H_{\rm eff}$, altogether resulting in a power-law decay of correlations also in the $k=0^{+}$ sector. Another, perhaps less likely scenario would instead be that the decays in both $k=0^{+}$ and $k=0^{-}$ stay as they are, an exponential and a power-law, respectively. In such a case the decay in the $k=0^{+}$ sector would eventually become faster than the one in $k=0^{-}$ (exponential decay always wins over a power law), meaning that the resonance in $k=0^{+}$ is not always the largest. Note that due to prefactors ($0.25\cdot 0.999986^{t}$ vs. $150/t^{3}$) $C(t)$ for the current $J$ would be larger than the one for $Z$ only for $t>2.5\cdot 10^{6}$ when $C(t)<10^{-17}$ (requiring $L>120$ to see it in a single random state).
@@
We find it quite interesting that one can extract all prethermalization information from the truncated operator propagator $M(k)$ – a procedure that seems simpler than for instance tediously evaluating BCH terms to a high order. It would be also interesting to understand physical mechanism behind this slow relaxation and whether behavior in different momentum sectors are somehow related. The slow breaking of integrability that we observes is a bit akin to what one expects in classical systems based on the KAM theorem. Upon weak perturbation some classical orbits (KAM torii) remain essentially integrable, e.g., display non-decaying autocorrelation functions, however in a high dimensional system they do not separate the phase space into inside and outside and orbits can escape (Arnold diffusion).
Refer to captionRefer to caption
Figure 12: An almost conserved operator in the mixing regime in sector $k=0^{+}$ for $\tau=0.45,\,h_{\rm x}=0.9,h_{\rm z}=0.8$. Top: Zoom of the spectrum around the largest $\lambda_{1}\approx 1-14\cdot 10^{-6}$ highlighted in red (blue circles are theoretical $r_{\rm in}$ and $r_{\rm out}$). Left is for $r=5$ (423 eigenvalues), right for $r=7$ (6303 eigenvalues). Bottom: The eigenvector $\mathbf{x}_{1}$ corresponding to $\lambda_{1}$ is for small $r$ equal to $H_{\rm eff}$ obtained by summing first $p$ BCH orders. Norm of the BCH term $H_{p}$ of order $p$ is shown with red diamonds.
@@ Appendix A Different parametrizations
where the strength of the field is given by $\cos{\gamma}=\cos{\alpha}\cos{\beta}$, while its direction is $\mathbf{n}=(\sin{\alpha}\cos{\beta},\sin{\alpha}\sin{\beta},\cos{\alpha}\sin{\beta})/\sin{\gamma}$. Because the interaction is invariant to rotations around the z axis, we can rotate the propagator around the z axis in order to have the field on the RHS of Eq.(26) only in the x-z plane. For the above $\mathbf{n}$ this is achieved by a rotation $V={\rm e}^{-{\rm i}\beta\sigma^{\rm z}/2}$, resulting in
@@ Appendix B Matrices
Rotation around the x-axis given by $U_{\rm x}={\rm e}^{{\rm i}\tau\sigma^{\rm x}}$ induces via $a^{\prime}=U_{\rm x}^{\dagger}aU_{\rm x}$ the following transformation
@@
At $\tau=\pi/4$ it is a signed permutation matrix, i.e., a permutation matrix will all diagonal matrix elements being zero, while half of off-diagonal ones is $-1$ (it is an antisymmetric matrix).

### v3

@@ IV.1 Power-law decay
Figure 10: Eigenvalue spectrum of $M(k=0^{-})$ for $\tau=0.45,h_{\rm x}=0.9,h_{\rm z}=0.8$ and $r=7$. There are $5985$ eigenvalues, with the largest at $\lambda_{1}\approx 0.914$ being highlighted by a red circle, with further 10 with the largest modulus being shown with small blue triangles (three largest ones are $\lambda_{2}\approx 0.897$, and a complex pair $|\lambda_{3,4}|\approx 0.861$). Blue circles are theoretical $r_{\rm in}$ and $r_{\rm out}$ (18).
@@ IV.2 Prethermalization in $k=0^{+}$
Considering the above explanation we must also rethink the apparent plateau seen in the $k=0^{+}$ sector (Fig. 8). Namely, it seems more natural that one would have a decay in all sectors. And indeed, looking at the autocorrelation function of $Z$ (20) at much longer times we eventually do see a decay, Fig. 11. At a finite $L$ there appears to be a 2nd plateau at a late time, e.g. $t>10^{5}$ at $L=20$, which though, as seen in the figure, is just a finite-size effect. As one increases $L$ one approaches the asymptotic exponential decay given by the largest eigenvalue that is very close to $1$, namely $\lambda_{1}\approx 1-14\cdot 10^{-6}$ (for studied truncations $\lambda_{1}$ seems rather stable with only small even-odd $r$ oscillations). This exponential decay only starts after very long transient 1st plateau of duration $\sim 10^{3}$. Looking at the $\tau$ dependence in the inset of Fig. 11 we see that the time and the height of this 2nd finite-size plateau rapidly change with $\tau$. We therefore have a rather interesting situation: observables from the $k=0^{-}$ sector decay (quickly) as a power law, while those from the $k=0^{+}$ eventually do decay exponentially but only after a long plateau.
Refer to captionRefer to caption
Figure 12: An almost conserved operator in the mixing regime in sector $k=0^{+}$ for $\tau=0.45,\,h_{\rm x}=0.9,h_{\rm z}=0.8$. Top: Zoom-in on the spectrum around the largest $\lambda_{1}\approx 1-14\cdot 10^{-6}$ highlighted in red (blue circles are theoretical $r_{\rm in}$ and $r_{\rm out}$), for $r=5$ (left, 423 eigenvalues), and $r=7$ (right, 6303 eigenvalues). Bottom: The eigenvector $\mathbf{x}_{1}$ corresponding to $\lambda_{1}$ is for small $r$ equal to $H_{\rm eff}$ obtained by summing first $p$ BCH orders. Norm of the BCH term $H_{p}$ of order $p$ is shown with red diamonds.
The long plateau can be explained by what has been called a prethermalization keiji16 . Namely, although $\tau$ does not necessarily look perturbatively small one can nevertheless try to write the effective one-step Floquet generator $H_{\rm eff}$ using the Baker-Campbell-Hausdorff (BCH) expansion
@@
with $H_{p}$ being the antihermitian $p$-th order term (in $\tau$) in the BCH expansion. If this expansion would converge, and therefore could describe the long-time dynamics by simple $U^{t}={\rm e}^{H_{\rm eff}t}$, the $H_{\rm eff}$ would be a constant of motion. Provided it would be a local operator this should be reflected in the spectral properties of $M(k=0^{+})$, i.e., one eigenvalue being $1$. However, the formal BCH expansion is usually divergent with terms $H_{p}$ eventually growing for $p>p_{0}$. Up until times when divergent $H_{p>p_{0}}$ become important the $H_{\rm eff}$ truncated at the order $p_{0}$ is almost conserved keiji16 , i.e., its exact lack of conservaton will become noticeable only at large times. Such an almost conserved $H_{\rm eff}$, a hallmark of prethermalization keiji16 , will result in one eigenvalue of $M(k=0^{+})$ being very close to $1$. And this is exactly what happens in our kicked Ising model at $\tau=0.45$. This can be seen in Fig. 12. As already mentioned, the largest eigenvalue $\lambda_{1}$ is very close to $1$ with the corresponding eigenvector $\mathbf{x}_{1}$ being, upto normalization, exactly equal to the truncated $H_{\rm eff}$. This is verified in the bottom frame of Fig. 12 where we calculate the overlaps between $\mathbf{x}_{1}$ calculated from $M(0^{+})$ for a finite support size $r$, say $r=3$, and compare it with $H_{\rm eff}$ obtained by summing up the first $p$ orders of the BCH series. Due to the structure of the kicked Ising model the support of the BCH term $H_{p}$ of order $p$ is $\sim 1+p/2$, i.e., every 2nd order increases the support of commutators by $1$ site. This can be nicely seen from the data: for $r=3$ the fidelity saturates at $p\approx 6$, while for $r=5$ at around $p\approx 10$. One can also see that the minimal error of $\mathbf{x}_{1}$ for a given $r$, i.e., its difference from $H_{\rm eff}$, is essentially equal to the norm $|H_{p}|^{2}:={\rm tr}(H_{p}^{\dagger}H_{p})$ of the truncated terms (red diamonds in the figure) for $p\approx 2r$, provided we are in a convergent regime $p<p_{0}$, that is, as long as $H_{p}$ decrease. Namely, by calculating BCH terms upto order $20$ casas08 we can see (Fig. 12) that the terms indeed start to increase after $p_{0}\approx 16$. This means that $H_{\rm eff}$ is not really conserved and we do not expect any eigenvalue of $M(k=0^{+})$ to be exactly $1$. The eventual divergent nature of the BCH series is also reflected in the structure of the eigenvector $\mathbf{x}_{1}$ – for larger $r$ it ceases to be well approximated by $H_{\rm eff}$ (compare $r=5$ and $r=7$ in the figure) – as well as in the spectrum of $M(0^{+})$ (other eigenvalues moving closer to $\lambda_{1}$).
@@
One question that remains is what happens in the thermodynamic limit $L\to\infty$, in particular in view of a seeming power-law decay in other sectors? There are two possibilities. As we can see in the top Fig. 12 there are many eigenvalues on the real axis, the number of which increases with $r$, similarly as is the case in the $k=0^{-}$ sector. So it could be that at very large $r$ when those eigenvalues move really close to $1$ they begin to mix with $\lambda_{1}$ which eventually loses its “prethermalization” character of describing $H_{\rm eff}$, altogether resulting in a power-law decay of correlations also in the $k=0^{+}$ sector. Another, perhaps less likely scenario would instead be that the decays in both $k=0^{+}$ and $k=0^{-}$ stay as they are in figures, an exponential and a power-law, respectively. In such a case the decay in the $k=0^{+}$ sector would eventually become faster than the one in $k=0^{-}$ (exponential decay always wins over a power law), meaning that the resonance in $k=0^{+}$ is not always the largest. Note that due to prefactors ($0.25\cdot 0.999986^{t}$ vs. $150/t^{3}$) $C(t)$ for the current $J$ would be larger than the one for $Z$ only for $t>2.5\cdot 10^{6}$ when $C(t)<10^{-17}$ (requiring $L>120$ to see it in a single random state).
@@
We find it quite interesting that one can extract all prethermalization information from the truncated operator propagator $M(k)$ – a procedure that seems simpler than for instance tediously evaluating BCH terms to a high order. It would be also interesting to understand the physical mechanism behind this slow relaxation and whether behavior in different momentum sectors are somehow related. A power-law decay of correlations at a finite $\tau=0.45$ means that in the thermodynamic limit and for finite breaking of integrability one can have situations without chaos with exponential decay of correlations. Such slow breaking of integrability is a bit akin to what one expects in classical systems based on the KAM theorem ott . Upon weak perturbation some classical orbits (KAM torii) remain essentially integrable, e.g., they display non-decaying autocorrelation functions. However, in a high dimensional system they do not separate the phase space into an inside and outside and orbits can escape (Arnold diffusion).
@@ Appendix A Different parametrizations
where the strength of the field is given by $\cos{\gamma}=\cos{\alpha}\cos{\beta}$, while its direction is $\mathbf{n}=(\sin{\alpha}\cos{\beta},\sin{\alpha}\sin{\beta},\cos{\alpha}\sin{\beta})/\sin{\gamma}$. Because the interaction is invariant to rotations around the z axis, we can rotate the propagator around the z axis in order to have the field on the RHS of Eq.(26) only in the x-z plane. For the above $\mathbf{n}$ this is achieved by rotation $V={\rm e}^{-{\rm i}\beta\sigma^{\rm z}/2}$, resulting in
@@ Appendix B Matrices
The rotation around the x-axis given by $U_{\rm x}={\rm e}^{{\rm i}\tau\sigma^{\rm x}}$ induces via $a^{\prime}=U_{\rm x}^{\dagger}aU_{\rm x}$ the following transformation
@@
At $\tau=\pi/4$ it is a signed permutation matrix, i.e., a permutation matrix with all diagonal matrix elements being zero, while half of the off-diagonal ones is $-1$ (it is an antisymmetric matrix).